OpenStudy (anonymous):

Integrate x^3 * sqrt(1-x^2).

OpenStudy (anonymous):

\[x^3(1-x^2)^{1/2}\] use u sub u= x^2 du= 2x dx

OpenStudy (anonymous):

sry! gotta go!